# A relationship between the Kauffman bracket skein algebras and Roger-Yang skein algebras of some small surfaces

### v1

@@ 1 Introduction
It is conjectured that the representation theory of the Roger-Yang skein algebra should encode hyperbolic geometric information about the punctured surface. In the analogous situation of the Kauffman bracket skein algebra for a closed surface, there is a Zariski dense subset where the maximal ideals of the center are in 1-1 correspondence with a point in the $\mathrm{SL}_{2}(\mathbb{C})$-character variety [20, 21, 24], and there are general constructions of irreducible, finite dimensional representations from geometric data [10, 11]. In addition, there are explicit descriptions of representations for the Kauffman bracket skein algebra for some small surfaces [36, 41]. In contrast, very little is known about the representation theory of the Roger-Yang skein algebra when the surface has at least two interior punctures (and no marked points on the boundary); see [25] for further discussion and references.
The goal of this paper is to examine one of the simpler non-trivial Roger-Yang skein algebra, namely that of the twice-punctured annulus. Let us denote it by $\mathcal{S}^{RY}(\Sigma_{0,2,2})$. We first determine a finite presentation for $\mathcal{S}^{RY}(\Sigma_{0,2,2})$; see Section 4. Note that presentations are known only for a few cases for the usual Kauffman bracket skein algebra [15, 17, 16] and the Roger-Yang skein algebra [7, 1].
@@
In Section 6, we characterize the representation theory of $\mathcal{S}^{RY}(\Sigma_{0,2,2})$ by adapting techniques for the closed torus from [36]. In particular, we show that when $N$ is a primitive root of unity with $N$ odd, then the central character of an irreducible, finite dimensional representation can be described by a complex 5-tuple, which we refer to as its classical shadow data. We show that if the 5-tuples satisfy certain polynomial conditions then the corresponding representations of $\mathcal{S}^{RY}(\Sigma_{0,2,2})$ are uniquely determined by their classical shadow data.
@@ 2 Acknowledgements
3 Definitions and Background
@@
We usually describe framed links using a diagram, which is a projection of the framed link that is isotoped into general position so that there are only transversal double points except at the punctures and so that the framing is always vertical. Over- and under-crossings at the double points are indicated by breaks in the projection. If there are more than two ends of arcs at a puncture, they are further labelled to show the ordering by height. Regular isotopy of the framed links in $\Sigma$ can be described using certain moves on their diagrams, as depicted in Figure 1.
@@
Let $R$ be a commutative domain, and $A\in R$ be an invertible element with a distinguished square roots $A^{\pm 1/2}$. For every $i=1,\ldots p$, we identify the $i$th puncture $v_{i}$ with an indeterminate variable.
@@
A simple framed link in $\Sigma\times[0,1]$ is one whose diagram contains no crossings and does not contain any loop bounding a disk or a disk containing exactly one puncture. The simple framed links freely span $\mathcal{S}^{RY}(\Sigma)$ as an $R$-module, [35].
Observe that the boundary components $\partial_{0},\ldots,\partial_{b}$ are central. When $A$ is a root of unity, there are additional central elements from threading by Chebyshev polynomials [25]. Let $T_{k}$ be the Chebyshev polynomial of the first kind, defined recursively by $T_{0}(x)=2$, $T_{1}(x)=x$, and $T_{k+1}(x)=xT_{k}(x)-T_{k-1}(x)$.
@@
For $A$ a primitive root of unity of odd order $N$, the center of the Roger-Yang skein algebra $Z(\mathcal{S}_{A}^{RY}(\Sigma))$ is the $\mathbb{C}[v_{i}^{\pm}]$-subalgebra generated by the following elements.
@@ 4 Presentations of $\mathcal{S}^{RY}(\Sigma_{0,2,2})$
The first presentation we provide involves the two boundary loops as well as the four generators shown in Figure 2. The relations from Theorem 4.1 are chosen for ease of proof rather than brevity. A more compact version will be presented later as Corollary 4.4.
@@
$0$ $=[\partial_{0},\partial_{1}]=[\partial_{0},\alpha]=[\partial_{0},\beta]=[\partial_{0},\gamma_{1}]=[\partial_{0},\gamma_{2}]=[\partial_{1},\alpha]=[\partial_{1},\beta]=[\partial_{1},\gamma_{1}]=[\partial_{1},\gamma_{2}]$ (8)
@@
Define $\overline{R}=\mathbb{Z}[A^{\pm 1/2},v_{i}^{\pm 1},\partial_{0}]$. Consider the alphabet $X=\{\alpha,\beta,\gamma_{1},\gamma_{2}\}$, and let $\langle X\rangle$ be the $\overline{R}$-algebra of finite words in this alphabet. Let $I$ be the ideal in $\langle X\rangle$ generated by relations (1)-(7).
@@
To use the Diamond Lemma, we construct a locally confluent terminating reduction system on $\langle X\rangle$. For definitions and more details about this approach, see [3]. For details relevant specifically to skein algebras, see [16, Section 9].
Define a reduction system $S$ on $\overline{R}\langle X\rangle$ by relations (1)-(7). Notice all the reduction rules are pairwise, meaning the left hand side is always the product of two letters in $X$. This implies that, to show this system is locally confluent, it is sufficient to show all overlap ambiguities are resolvable. Here, an overlap ambiguity is $x_{1}x_{2}x_{3}\in\langle X\rangle$ such that $x_{1}x_{2}$ and $x_{2}x_{3}$ both have reduction rules. It is resolvable if there exist sequences of reductions to $x_{1}x_{2}x_{3}$, beginning with the rules for $x_{1}x_{2}$ and $x_{2}x_{3}$ respectively that agree at their final expression. There are finitely many overlap ambiguities, so this can be checked by straightforward (if tedious) calculations.
@@
We now give an explicit description of this basis. Let $m=y_{1}\cdots y_{k}$ be an irreducible monomial and order $X$ by $\alpha<\beta<\gamma_{1}<\gamma_{2}$. For each pair $(x_{1},x_{2})$ with $x_{1}>x_{2}$ there is a reduction rule, so we must have $y_{i}\leq y_{i+1}$ for each $i$. Hence, $m=\alpha^{e_{1}}\beta^{e_{2}}\gamma_{1}^{e_{3}}\gamma_{2}^{e_{4}}$ for some $e_{i}\in\mathbb{Z}_{\geq 0}$. We can apply (6) to further reduce the monomial if and only if $e_{3}e_{4}\neq 0$. Thus, the set $B=\{\alpha^{e_{1}}\beta^{e_{2}}\gamma_{1}^{e_{3}}\gamma_{2}^{e_{4}}\mid e_{3}e_{4}=0\}$ is an $\overline{R}$-basis for $\overline{R}\langle X\rangle/I$.
@@
To upgrade $\psi$ to an isomorphism, it remains to show $\psi(B)$ is $\overline{R}$-linearly independent in $\mathcal{S}^{RY}(\Sigma_{0,2,2})$. Our strategy is to quotient $\mathcal{S}^{RY}(\Sigma_{0,2,2})$ even further, and show the image of $\psi(B)$ under this quotient mapping is linearly independent. We use the following reformulation of [15, Lemma 1.2].
@@
In our case, $S$ is $\mathcal{S}^{RY}(\Sigma_{0,2,2})$, which is torsion free by [35, Theorem 2.4] and $\mathcal{R}$ is $\overline{R}$, which is clearly Noetherian. The map $\pi$ is given by the following chain of maps.
@@
We end this section by providing a simpler presentation for $\mathcal{S}^{RY}(\Sigma_{0,2,2})$.
Definition 4.3.
Given $\alpha,\beta\in\mathcal{S}^{RY}(\Sigma)$, define $\alpha,\beta]_{A}=A\alpha*\beta-A^{-1}\beta*\alpha$.
@@ 5 Relationships between skein algebras of small surfaces
is an $\overline{R}$-basis for $\mathcal{S}^{RY}(\Sigma_{0,2,2})$. The image of this basis under $\varphi$ is the basis for $\mathcal{S}(\Sigma_{1,0,0})$ constructed in [15, Thm. 2.1]. Hence, $\varphi$ is an extension of a ring homomorphism $\overline{R}\to\mathbb{Z}[A^{\pm 1/2}]$, so its kernel is the ideal generated in $\mathcal{S}^{RY}(\Sigma_{0,2,2})$ by some ideal of $\overline{R}\cdot\emptyset$ (recall $\emptyset$ denotes the empty link), and $\overline{R}\cdot\emptyset$ is the center of $\mathcal{S}^{RY}(\Sigma_{0,2,2})$. ∎
@@
There is a surjective homomorphism $\mathcal{S}^{RY}(\Sigma_{0,1,3})\to\mathcal{S}(\Sigma_{1,1,0})$ given by identification of generators and sending $A\mapsto A^{2}$ and $v_{i}\mapsto 1$. Hence, $\mathcal{S}(\Sigma_{1,1,0})\cong\mathcal{S}^{RY}(\Sigma_{0,1,3})/(v_{i}-1)$.
@@
In [1, Theorem 1.1], the authors prove that $\mathcal{S}^{RY}(\Sigma_{0,1,3})$ is the $\mathbb{Z}[A^{\pm 1/2},v_{i}^{\pm 1}]$-algebra generated by $x_{1},x_{2},x_{3}$ subject to the relation $v_{i}[x_{i},x_{i+1}]_{A^{1/2}}=(A-A^{-1})x_{i+2}$ with subscripts taken mod 3. In [15, Theorem 2.1], the authors show $\mathcal{S}(\Sigma_{1,1,0})$ is the $\mathbb{Z}[A^{\pm 1}]$-algebra generated by $x_{1},x_{2},x_{3}$ subject to the relation $x_{i},x_{i+1}]_{A}=(A^{2}-A^{-2})x_{i+2}$ with subscripts taken mod 3. From this, it is clear that the map described in the statement is a homomorphism with kernel $v_{i}-1$. ∎
@@
6 Classification of Representations of $\mathcal{S}^{RY}(\Sigma_{0,2,2})$
In this section, we study the representations of $\mathcal{S}^{RY}(\Sigma_{0,2,2})$ where $A$ is a root of unity. In particular, we construct representations of $\mathcal{S}^{RY}(\Sigma_{0,2,2})$ from the data of its classical shadow.
Throughout this section, let $A$ be a primitive $N$th root of 1-1, with $N$ odd. Let $\rho$ be an irreducible representation of $\mathcal{S}^{RY}(\Sigma_{0,2,2})$ over a $\mathbb{C}$-vector space $V$. By Schur’s Lemma, we can associate a scalar to each element of the center $Z(\mathcal{S}^{RY}(\Sigma_{0,2,2}))$. Let $\chi_{\rho}:Z(\mathcal{S}^{RY}(\Sigma_{0,2,2}))\to\mathbb{C}$ denote the corresponding central character. By Proposition 3.3 and the presentation of Corollary 4.4, $\chi_{\rho}$ is determined entirely by the values it takes at the generators $\alpha,\beta,\gamma,\partial_{0}$, and $\partial_{1}$, namely by the 5-tuple $(t_{1},t_{2},t_{3},d_{0},d_{1})\in\mathbb{C}^{5}$ defined by

### v2

@@ 1 Introduction
It is conjectured that the representation theory of the Roger-Yang skein algebra should encode hyperbolic geometric information about the punctured surface. In the analogous situation of the Kauffman bracket skein algebra for a surface, there is a Unicity Theorem that says that a Zariski open dense subset of the $\mathrm{SL}_{2}(\mathbb{C})$-character variety of the surface where points in the subset are in one-to-one correspondence with an irreducible, finite-dimensional representation of the Kauffman bracket skein algebra [23, 24, 27]. Moreover, there are general constructions of irreducible, finite dimensional representations from geometric data [12, 13], and for some small surfaces, these representations have been explicitly computed [41, 46]. In contrast, much less is known about the representation theory of the Roger-Yang skein algebra when the surface has at least two interior punctures (and no marked points on the boundary); see [28] for further discussion and references.
The goal of this paper is to examine one of the simpler non-trivial Roger-Yang skein algebra, namely that of the twice-punctured annulus. Let us denote it by $\mathcal{S}^{RY}(\Sigma_{0,2,2})$. We first determine a finite presentation for $\mathcal{S}^{RY}(\Sigma_{0,2,2})$ where the generators can be chosen to be ordered (as in [1]); see Section 4. Note that presentations are known only for a few cases for the usual Kauffman bracket skein algebra [17, 20, 19] and the Roger-Yang skein algebra [8, 2].
@@
In Section 6, we study the representation theory of $\mathcal{S}^{RY}(\Sigma_{0,2,2})$ following the work of [23, 41, 26] for the closed torus and $U_{q}(so_{3})$. We first show that, like the usual Kauffman bracket skein algebra, $\mathcal{S}^{RY}(\Sigma_{0,2,2})$ is almost Azumaya, which implies that almost every irreducible, finite-dimensional representation of $\mathcal{S}^{RY}(\Sigma_{0,2,2})$ is determined by its central character. When $A$ is a primitive $2N$ root of unity with $N$ odd, the central character can be described by a complex 5-tuple, which we refer to as its classical shadow data. We provide polynomial conditions under which the classical shadow data uniquely determines a representation. We also provide explicit computations of the representations, which are all $N$-dimensional.
@@ 2 Acknowledgements
3 Definition of Roger-Yang skein algebra
@@
We usually describe framed links using a diagram, that is, a projection of the framed link which is isotoped into general position. This means there are only transversal double points except at the punctures and the framing is always vertical. Over- and under-crossings at the double points are indicated by breaks in the projection. If there are more than two ends of arcs at a puncture, they are further labelled to show the ordering by height. Regular isotopy of the framed links in $\Sigma$ can be described using certain moves on their diagrams, as depicted in Figure 1.
@@
Let $R$ be a commutative domain, and $A\in R$ be an invertible element with distinguished square roots $A^{\pm 1/2}$. For every $i=1,\ldots p$, we identify the $i$th puncture with the indeterminate variable $v_{i}$.
@@
Note that, in contrast, the usual Kauffman bracket skein algebra of $\mathcal{S}(\Sigma)$ does not include vertex classes or framed arcs. More specifically, $\mathcal{S}(\Sigma)$ is the $R$-algebra freely generated by disjoint unions of framed knots in $\Sigma\times[0,1]$ modded out by relations 1) and 3) from Definition 3.1.
We next state some results about the structure of $\mathcal{S}^{RY}(\Sigma)$ that will be used in the paper.
We say that a framed link in $\Sigma\times[0,1]$ is simple if it admits a diagram that contains no crossings and does not contain any loop bounding a disk with one or no punctures. The empty link, denoted $\emptyset$, is considered a simple link.
@@
The simple framed links freely span $\mathcal{S}^{RY}(\Sigma)$ as an $R$-module.
Proposition 3.4 ( [7] ).
$\mathcal{S}^{RY}(\Sigma)$ is orderly finitely generated as an algebra, meaning that there exists skein elements $a_{1},\dots,a_{k}$ such that the set of all ordered monomials $\{a_{1}^{n_{1}}\cdots a_{k}^{n_{k}}\}_{n_{i}\in\mathbb{Z}_{\geq 0}}$ spans $\mathcal{S}^{RY}(\Sigma)$.
It is an open question whether there is an orderly generating set consisting of simple curves. In particular, it is not known whether the orderly generating set from [7] are simple curves. On the other hand, there is a generating set from [9] consisting of simple curves, but it is not known whether they form an orderly finitely generating set for most surfaces. In Section 4, we will show that $\mathcal{S}^{RY}(\Sigma_{0,2,2})$ is orderly generated using simple curves.
Proposition 3.5 ( [35, 7] ).
$\mathcal{S}^{RY}(\Sigma)$ is an integral domain.
Let $T_{k}$ be the Chebyshev polynomial of the first kind, defined recursively by $T_{0}(x)=2$, $T_{1}(x)=x$, and $T_{k+1}(x)=xT_{k}(x)-T_{k-1}(x)$.
Proposition 3.6 ([28], following [11, 32, 29]).
If $A^{2N}=1$ with $N$ odd, the center of the Roger-Yang skein algebra $Z(\mathcal{S}_{A}^{RY}(\Sigma))$ contains the following skeins:
@@
Moreover, if $A$ is a primitive root of unity of odd order $N$, then the center $Z(\mathcal{S}_{A}^{RY}(\Sigma))$ is equal to the $\mathbb{C}[v_{i}^{\pm}]$-subalgebra generated by the above skeins.
Finally, we introduce some notation.
Definition 3.7.
Given $\alpha,\beta\in\mathcal{S}^{RY}(\Sigma)$, define $\alpha,\beta]=\alpha*\beta-\beta*\alpha$ and $\alpha,\beta]_{A}=A\alpha*\beta-A^{-1}\beta*\alpha$.
@@ 4 Presentations of $\mathcal{S}^{RY}(\Sigma_{0,2,2})$
The first presentation we provide involves one of the boundary loops as well as the four generators shown in Figure 2. The relations from Theorem 4.1 are chosen for ease of proof rather than brevity. A more compact version will be presented later as Corollary 4.4.
@@
$0$ $=[\partial_{0},\partial_{1}]=[\partial_{0},\alpha]=[\partial_{0},\beta]=[\partial_{0},\gamma_{1}]=[\partial_{0},\gamma_{2}]$ (8)
@@
Define $\overline{R}=\mathbb{Z}[A^{\pm 1/2},v_{i}^{\pm 1},\partial_{0}]$. Consider the alphabet $X=\{\alpha,\beta,\gamma_{1},\gamma_{2}\}$, and let $\langle X\rangle$ be the $\overline{R}$-algebra of finite words in this alphabet. Let $I$ be the $\overline{R}$-ideal in $\langle X\rangle$ generated by relations (1)-(7).
@@
To use the Diamond Lemma, we construct a locally confluent terminating reduction system on $\langle X\rangle$. For definitions and more details about this approach, see [4]. For details relevant specifically to skein algebras, see [19, Section 9].
Define a reduction system $S$ on $\overline{R}\langle X\rangle$ by relations (1)-(7). Notice all the reduction rules are pairwise, meaning the left hand side is always the product of two letters in $X$. This implies that, to show this system is locally confluent, it is sufficient to show all overlap ambiguities are resolvable. Here, an overlap ambiguity is $x_{1}x_{2}x_{3}\in\langle X\rangle$ such that $x_{1}x_{2}$ and $x_{2}x_{3}$ both have reduction rules. It is resolvable if there exist sequences of reductions of $x_{1}x_{2}x_{3}$, beginning with the rules for $x_{1}x_{2}$ and $x_{2}x_{3}$ respectively that agree at their final expression. There are finitely many overlap ambiguities, so this can be checked by straightforward (if tedious) calculations.
@@
We now give an explicit description of this basis. Let $m=x_{i_{1}}\cdots x_{i_{k}}$ be an irreducible monomial and recall there is an order on $X$ given by $\alpha<\beta<\gamma_{1}<\gamma_{2}$. For each pair $(x_{1},x_{2})$ with $x_{1}>x_{2}$ there is a reduction rule, so we must have $y_{i_{j}}\leq y_{i_{j+1}}$ for each $j$. Hence, $m=\alpha^{e_{1}}\beta^{e_{2}}\gamma_{1}^{e_{3}}\gamma_{2}^{e_{4}}$ for some $e_{i}\in\mathbb{Z}_{\geq 0}$. We can apply (6) to further reduce the monomial if and only if $e_{3}e_{4}\neq 0$. Thus, the set $B=\{\alpha^{e_{1}}\beta^{e_{2}}\gamma_{1}^{e_{3}}\gamma_{2}^{e_{4}}\mid e_{3}e_{4}=0\}$ is an $\overline{R}$-basis for $\overline{R}\langle X\rangle/I$.
@@
To upgrade $\psi$ to an isomorphism, it remains to show $\psi(B)$ is $\overline{R}$-linearly independent in $\mathcal{S}^{RY}(\Sigma_{0,2,2})$. Our strategy is to quotient $\mathcal{S}^{RY}(\Sigma_{0,2,2})$ even further, and show the image of $\psi(B)$ under this quotient mapping is linearly independent. We use the following reformulation of [17, Lemma 1.2].
@@
In our case, $S$ is $\mathcal{S}^{RY}(\Sigma_{0,2,2})$, which is torsion free by [40, Theorem 2.4] and $\mathcal{R}$ is $\overline{R}$, which is clearly Noetherian. The map $\pi$ is given by the following chain of maps.
@@
Corollary 4.3.
$\mathcal{S}^{RY}(\Sigma_{0,2,2})$ is orderly finitely generated by simple curves.
Proof.
It is shown in the proof of Theorem 4.1 that $B=\{\alpha^{e_{1}}\beta^{e_{2}}\gamma_{1}^{e_{3}}\gamma_{2}^{e_{4}}\mid e_{3}e_{4}=0\}$ is a basis for $\mathcal{S}^{RY}(\Sigma_{0,2,2})$. So the simple curves $\alpha,\beta,\gamma_{1},\gamma_{2}$ form an orderly finitely generating set. ∎
We end this section by providing a simpler presentation for $\mathcal{S}^{RY}(\Sigma_{0,2,2})$. We use the notation from Definition 3.7: $\alpha,\beta]_{A}=A\alpha*\beta-A^{-1}\beta*\alpha$.
@@ 5 Relationships between skein algebras of small surfaces
is an $\overline{R}$-basis for $\mathcal{S}^{RY}(\Sigma_{0,2,2})$. The image of this basis under $\varphi$ is the basis for $\mathcal{S}(\Sigma_{1,0,0})$ constructed in [17, Thm. 2.1]. Hence, $\varphi$ is an extension of a ring homomorphism $\overline{R}\to\mathbb{Z}[A^{\pm 1/2}]$, so its kernel is the ideal generated in $\mathcal{S}^{RY}(\Sigma_{0,2,2})$ by some ideal of $\overline{R}\cdot\emptyset$ (recall $\emptyset$ denotes the empty link), and $\overline{R}\cdot\emptyset$ is the center of $\mathcal{S}^{RY}(\Sigma_{0,2,2})$. ∎
@@
There is a surjective homomorphism $\mathcal{S}^{RY}(\Sigma_{0,1,3})\to\mathcal{S}(\Sigma_{1,1,0})$ given by identification of generators and sending $A^{1/2}\mapsto A$ and $v_{i}\mapsto 1$. This induces an isomorphism, $\mathcal{S}(\Sigma_{1,1,0})\cong\mathcal{S}^{RY}(\Sigma_{0,1,3})/(v_{i}-1)$.
@@
In [2, Theorem 1.1], the authors prove that $\mathcal{S}^{RY}(\Sigma_{0,1,3})$ is the $\mathbb{Z}[A^{\pm 1/2},v_{i}^{\pm 1}]$-algebra generated by $x_{1},x_{2},x_{3}$ subject to the relation $v_{i}[x_{i},x_{i+1}]_{A^{1/2}}=(A-A^{-1})x_{i+2}$ with subscripts taken mod 3. In [17, Theorem 2.1], the authors show $\mathcal{S}(\Sigma_{1,1,0})$ is the $\mathbb{Z}[A^{\pm 1}]$-algebra generated by $x_{1},x_{2},x_{3}$ subject to the relation $x_{i},x_{i+1}]_{A}=(A^{2}-A^{-2})x_{i+2}$ with subscripts taken mod 3. From this, it is clear that the map described in the statement is a homomorphism with kernel $v_{i}-1$. ∎
@@
6 Representation theory of $\mathcal{S}^{RY}(\Sigma_{0,2,2})$
In this section, we consider $\mathcal{S}^{RY}(\Sigma)$ as a $\mathbb{C}$-algebra instead of a $\mathbb{Z}$-algebra.
6.1 $\mathcal{S}^{RY}(\Sigma_{0,2,2})$ is almost Azumaya
We refer the interested reader to [15, 23] for definitions and details about almost Azumaya algebras.
For $\mathcal{S}^{RY}(\Sigma_{0,2,2})$, the following proof was suggested by Thang Lê (see Remark 5.6 in [28]).
Theorem 6.1.
When $A^{2N}=1$, $\mathcal{S}^{RY}(\Sigma_{0,2,2})$ is almost Azumaya.
Proof.
To show that an algebra is almost Azumaya, we use Theorem 3.6 from [23], which shows an algebra is almost Azumaya if it satisifes three conditions: it is finitely generated as a $\mathbb{C}$-algebra, it is prime, and it is finitely generated over its center.
For $\mathcal{S}^{RY}(\Sigma_{0,2,2})$, the first two conditions follow from [9, 35, 7]. To prove that $\mathcal{S}^{RY}(\Sigma_{0,2,2})$ is finitely generated over its center, recall that Corollary 4.3 shows $\mathcal{S}^{RY}(\Sigma_{0,2,2})$ is orderly finitely generated by simple loops and arcs $B=\{\alpha^{e_{1}}\beta^{e_{2}}\gamma_{1}^{e_{3}}\gamma_{2}^{e_{4}}\mid e_{3}e_{4}=0\}$. Thus application of the Chebyshev polynomial from Proposition 3.6 produces central elements $T_{N}(\alpha)$, $T_{N}(\beta)$, $T_{N}(\sqrt{v_{1}v_{2}}\gamma_{1})$, and $T_{N}(\sqrt{v_{1}v_{2}}\gamma_{2})$. Note that the vertex classes $v_{1}$ and $v_{2}$ are scalars and hence commutative. We can thus decrease the degree on $\alpha,\beta,\gamma_{1}$ and $\gamma_{2}$ until it is equal to or less than $N$, the degree of $T_{N}$. Thus $\mathcal{S}^{RY}(\Sigma_{0,2,2})$ is finite dimensional over its center, generated by $\{\alpha^{e_{1}}\beta^{e_{2}}\gamma_{1}^{e_{3}}\gamma_{2}^{e_{4}}\mid 0\leq e_{i}<N,\;e_{3}e_{4}=0\}$ as a $Z(\mathcal{S}^{RY}(\Sigma_{0,2,2}))$-module.
We thus have implications for the representation theory for $\mathcal{S}^{RY}(\Sigma_{0,2,2})$, analogous to the Unicity Theorem for the usual Kauffman bracket skein algebra. For an algebra $\mathcal{A}$, let $\mathrm{Irrep}(\mathcal{A})$ be the set of all equivalence classes of finite-dimensional irreducible representation of $\mathcal{A}$ and let $\mathrm{MaxSpec}(Z(\mathcal{A}))$ be the set of all maximal ideals of its center $Z(\mathcal{A})$. For every finite-dimensional irreducible representation $\rho$ of $\mathcal{A}$, one may define its associated central character by using Schur’s Lemma to associate a scalar to each element of the center $Z(A)$. Up to isomorphism, the central character is in bijection with the maximal ideals of $Z(\mathcal{A})$. Thus, there is a well-defined map $\chi:\mathrm{Irrep}(\mathcal{A})\to\mathrm{MaxSpec}\;(\mathcal{A})$.
Corollary 6.2.
When $\mathcal{A}=\mathcal{S}^{RY}(\Sigma_{0,2,2})$, and $D$ is the dimension of $\mathcal{A}$ as a $Z(\mathcal{A})$-module, every irreducible representation of $\mathcal{S}^{RY}(\Sigma_{0,2,2})$ has dimension at most $\sqrt{D}$, and $\chi:\mathrm{Irrep}(\mathcal{A})\to\mathrm{MaxSpec}\;Z(\mathcal{A})$ is surjective. Moreover, there exists a Zariski open dense subset $U\subset\mathrm{MaxSpec}\;Z(\mathcal{A})$ on which $\chi^{-1}$ is injective, and the dimension of any irreducible representation in $\chi^{-1}(U)$ is equal to $\sqrt{D}$.
In the next section, we provide explicit computations on the Zariski open dense subset $U$ and show that the associated representations have dimension $\sqrt{D}=N$ when $A^{2N}=1$ and $N$ is odd.
6.2 Explicit representations of $\mathcal{S}^{RY}(\Sigma_{0,2,2})$
Throughout this section, let $A$ be a primitive $2N$th root of unity with $N$ odd. Let $\rho$ be an irreducible representation of $\mathcal{S}^{RY}(\Sigma_{0,2,2})$ over a $\mathbb{C}$-vector space $V$. By Schur’s Lemma, we can associate a scalar to each element of the center $Z(\mathcal{S}^{RY}(\Sigma_{0,2,2}))$. Let $\chi_{\rho}:Z(\mathcal{S}^{RY}(\Sigma_{0,2,2}))\to\mathbb{C}$ denote the corresponding central character. We evaluate $\chi_{p}$ at central elements corresponding to the generators to obtain the 5-tuple $(t_{1},t_{2},t_{3},d_{0},d_{1})\in\mathbb{C}^{5}$ defined by